im trying to figure out how to get The width of a rectangular feed lot is 8 m less
than the length. The area is 20 m squared . Find the length and
the width.
i thought it would be (8-m)((20-m)squared

Answer by ankor@dixie-net.com(22740)

The width of a rectangular feed lot is 8 m less than the length.
The area is 20 m squared .
Find the length and the width.
:
"The width of a rectangular feed lot is 8 m less than the length."
W = (L-8)
:
L * W = 20
Replace W with (L-8), find L
L * (L-8) = 20
A quadratic equation
L^2 - 8L - 20 = 0
factors to
(L-10)*L+2) = 0
The positive solution is what we want here
L = 10 m is the length
then
W = 10 - 8
W = 2 m is the width